# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ Appendix C Interferometric BB84 With Added Measurements
In this subsection, we revisit the attack on the interferometric BB84 implementation from Appendix B. Recall that a Reversed-Space Attack stems from Eve being able to use time modes other than $t^{\prime}_{0}$ and $t^{\prime}_{1}$ in order to fool Bob’s detection device.
However, by doing so, Eve creates pulses at times $t_{3}$ and $t_{-1}$ at Bob’s device, while such modes can never happen if only the incoming qubits arrive in time mode $t^{\prime}_{0}$ and $t^{\prime}_{1}$, as they should be.
@@ C.1 Bob’s Measured Space
$\mathcal{H}^{\mathrm{B}}=\operatorname{span}\left\{|{V}\rangle,|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{d_{0}}\rangle,|{d_{0}}\rangle,|{s_{1}}\rangle,|{d_{1}}\rangle,|{s_{2}}\rangle,|{d_{2}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.$ (33)
Bob’s unitary is the same for both measurement bases, and is identical to the one given in Section 5.3 ($\phi=0$ still holds):
$U_{\mathrm{B}}|{t^{\prime}_{i}}\rangle=\left(|{s_{t}}\rangle+i|{d_{t}}\rangle-|{s_{t+1}}\rangle+i|{d_{t+1}}\rangle\right)/\thickspace 2.$ (34)
@@
$\begin{array}[]{rcl}J_{0}&=&\left\{|{s_{0}}\rangle,|{d_{0}}\rangle\right\},\\ J_{1}&=&\left\{|{s_{2}}\rangle,|{d_{2}}\rangle\right\},\\ J_{{\operatorname{loss}}}&=&\left\{|{d_{1}}\rangle,|{s_{1}}\rangle\right\},\\ J_{{\operatorname{invalid}}}&=&\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.\end{array}$ (35)
@@
$\begin{array}[]{rcl}J_{0}&=&\left\{|{d_{1}}\rangle\right\},\\ J_{1}&=&\left\{|{s_{1}}\rangle\right\},\\ J_{{\operatorname{loss}}}&=&\left\{|{s_{0}}\rangle,|{d_{0}}\rangle,|{s_{2}}\rangle,|{d_{2}}\rangle\right\},\\ J_{{\operatorname{invalid}}}&=&\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.\end{array}$ (36)
C.2 Eve’s Attack
@@
$\mathcal{H}^{\mathrm{P}}=\operatorname{span}\{|{t^{\prime}_{-2}}\rangle,|{t^{\prime}_{-1}}\rangle|{t^{\prime}_{0}}\rangle,|{t^{\prime}_{1}}\rangle,|{t^{\prime}_{2}}\rangle,|{t^{\prime}_{3}}\rangle\}.$ (37)
@@ C.3 Attack Constraints
$J_{{\operatorname{error}}}=J_{1}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{s_{2}}\rangle,|{s_{3}}\rangle,|{d_{-1}}\rangle,|{d_{2}}\rangle,|{d_{3}}\rangle\right\}.$ (39)
@@
    $\epsilon_{0,-2}|{E_{0,-2}}\rangle-\epsilon_{0,-1}|{E_{0,-1}}\rangle=0.$ (40)
@@
    $\epsilon_{0,1}|{E_{0,1}}\rangle-\epsilon_{0,2}|{E_{0,2}}\rangle=0.$ (41)
@@
    $\begin{array}[]{c}\epsilon_{0,2}|{E_{0,2}}\rangle-\epsilon_{0,3}|{E_{0,3}}\rangle=0\\ \Downarrow\\ \epsilon_{0,1}|{E_{0,1}}\rangle=\epsilon_{0,2}|{E_{0,2}}\rangle=\epsilon_{0,3}|{E_{0,3}}\rangle.\end{array}$ (42)
@@
    $\begin{array}[]{c}\epsilon_{0,-2}|{E_{0,-2}}\rangle+\epsilon_{0,-1}|{E_{0,-1}}\rangle=0\\ \Downarrow\\ \epsilon_{0,-2}=\epsilon_{0,-1}=0.\end{array}$ (43)
@@
    $\begin{array}[]{c}\epsilon_{0,1}|{E_{0,1}}\rangle+\epsilon_{0,2}|{E_{0,2}}\rangle=0\\ \Downarrow\\ \epsilon_{0,1}=\epsilon_{0,2}=\epsilon_{0,3}=0.\end{array}$ (44)
@@
    $\begin{array}[]{c}\underset{=0}{\epsilon_{0,2}}|{E_{0,2}}\rangle+\underset{=0}{\epsilon_{0,3}}|{E_{0,3}}\rangle=0.\end{array}$ (45)
@@
$J_{{\operatorname{error}}}=J_{0}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{s_{0}}\rangle,|{s_{3}}\rangle,|{d_{-1}}\rangle,|{d_{0}}\rangle,|{d_{3}}\rangle\right\}.$ (46)
@@
    $\epsilon_{1,-2}|{E_{1,-2}}\rangle-\epsilon_{1,-2}|{E_{1,-2}}\rangle=0.$ (47)
@@
    $\begin{array}[]{c}\epsilon{1,-1}|{E_{1,-1}}\rangle-\epsilon_{1,0}|{E_{1,0}}\rangle=0\\ \Downarrow\\ \epsilon_{1,-2}=\epsilon_{1,-1}=\epsilon_{1,0}.\end{array}$ (48)
@@
    $\epsilon_{1,2}|{E_{1,2}}\rangle-\epsilon_{1,3}|{E_{1,3}}\rangle=0.$ (49)
@@
    $\begin{array}[]{c}\epsilon_{1,-2}|{E_{1,-2}}\rangle+\epsilon_{1,-2}|{E_{1,-2}}\rangle=0\\ \Downarrow\\ \epsilon_{1,-2}=\epsilon_{1,-1}=\epsilon_{1,0}=0.\end{array}$ (50)
@@
    $\underset{=0}{\epsilon_{1,-1}}|{E_{1,-1}}\rangle+\underset{=0}{\epsilon_{1,0}}|{E_{1,0}}\rangle=0.$ (51)
@@
    $\begin{array}[]{c}\epsilon_{1,2}|{E_{1,2}}\rangle+\epsilon_{1,3}|{E_{1,3}}\rangle=0\\ \Downarrow\\ \epsilon_{1,2}=\epsilon_{1,3}=0.\end{array}$ (52)
@@
$J_{{\operatorname{error}}}=J_{1}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.$ (54)
@@
$\begin{array}[]{c}\sum_{k}\epsilon_{0,k}\beta_{k,d_{1}}|{E_{0,k}}\rangle+\sum_{k}\epsilon_{1,k}\beta_{k,d_{1}}|{E_{1,k}}\rangle=0\\ \Downarrow\\ \frac{i}{\sqrt{2}}(-\epsilon_{0,0}|{E_{0,0}}\rangle+\underset{=0}{\epsilon_{0,1}}|{E_{0,1}}\rangle)+\frac{i}{\sqrt{2}}(-\underset{=0}{\epsilon_{1,0}}|{E_{1,0}}\rangle+\epsilon_{1,1}|{E_{1,1}}\rangle)=0\\ \Downarrow\\ \epsilon_{0,0}|{E_{0,0}}\rangle=\epsilon_{1,1}|{E_{1,1}}\rangle.\end{array}$ (55)
@@
$\begin{array}[]{c}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}},\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}\\ \Downarrow\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{A}}.\end{array}$ (56)

### v3

@@ Appendix C Interferometric BB84 With Added Measurements
In this appendix, we revisit the attack on the interferometric BB84 implementation from Appendix B, in which Eve uses time modes other than $t^{\prime}_{0}$ and $t^{\prime}_{1}$ in order to fool Bob’s detection device.
However, by doing so, Eve creates pulses at times $t_{3}$ and $t_{-1}$ at Bob’s device, while such modes can never happen if only the incoming qubits arrive in time mode $t^{\prime}_{0}$ and $t^{\prime}_{1}$, as they should.
@@ C.1. Bob’s Measured Space
(33) $\mathcal{H}^{\mathrm{B}}=\operatorname{span}\left\{|{V}\rangle,|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{d_{0}}\rangle,|{d_{0}}\rangle,|{s_{1}}\rangle,|{d_{1}}\rangle,|{s_{2}}\rangle,|{d_{2}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.$
Bob’s unitary is the same for both measurement bases, and is identical to the one given in Eq. (15) ($\theta=0$ still holds):
(34) $U_{\mathrm{B}}|{t^{\prime}_{n}}\rangle=\left(|{s_{n}}\rangle+i|{d_{n}}\rangle-|{s_{n+1}}\rangle+i|{d_{n+1}}\rangle\right)/\thickspace 2.$
@@
(35) $\begin{array}[]{rcl}J_{0}&=&\left\{|{s_{0}}\rangle,|{d_{0}}\rangle\right\},\\ J_{1}&=&\left\{|{s_{2}}\rangle,|{d_{2}}\rangle\right\},\\ J_{{\operatorname{loss}}}&=&\left\{|{d_{1}}\rangle,|{s_{1}}\rangle\right\},\\ J_{{\operatorname{invalid}}}&=&\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.\end{array}$
@@
(36) $\begin{array}[]{rcl}J_{0}&=&\left\{|{d_{1}}\rangle\right\},\\ J_{1}&=&\left\{|{s_{1}}\rangle\right\},\\ J_{{\operatorname{loss}}}&=&\left\{|{s_{0}}\rangle,|{d_{0}}\rangle,|{s_{2}}\rangle,|{d_{2}}\rangle\right\},\\ J_{{\operatorname{invalid}}}&=&\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.\end{array}$
C.2. Eve’s Attack
@@
(37) $\mathcal{H}^{\mathrm{P}}=\operatorname{span}\{|{t^{\prime}_{-2}}\rangle,|{t^{\prime}_{-1}}\rangle|{t^{\prime}_{0}}\rangle,|{t^{\prime}_{1}}\rangle,|{t^{\prime}_{2}}\rangle,|{t^{\prime}_{3}}\rangle\}.$
@@ C.3. Attack Constraints
(39) $J_{{\operatorname{error}}}=J_{1}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{s_{2}}\rangle,|{s_{3}}\rangle,|{d_{-1}}\rangle,|{d_{2}}\rangle,|{d_{3}}\rangle\right\}.$
@@
    (40) $\epsilon_{0,-2}|{E_{0,-2}}\rangle-\epsilon_{0,-1}|{E_{0,-1}}\rangle=0.$
@@
    (41) $\epsilon_{0,1}|{E_{0,1}}\rangle-\epsilon_{0,2}|{E_{0,2}}\rangle=0.$
@@
    (42) $\begin{array}[]{c}\epsilon_{0,2}|{E_{0,2}}\rangle-\epsilon_{0,3}|{E_{0,3}}\rangle=0\\ \Downarrow\\ \epsilon_{0,1}|{E_{0,1}}\rangle=\epsilon_{0,2}|{E_{0,2}}\rangle=\epsilon_{0,3}|{E_{0,3}}\rangle.\end{array}$
@@
    (43) $\begin{array}[]{c}\epsilon_{0,-2}|{E_{0,-2}}\rangle+\epsilon_{0,-1}|{E_{0,-1}}\rangle=0\\ \Downarrow\\ \epsilon_{0,-2}=\epsilon_{0,-1}=0.\end{array}$
@@
    (44) $\begin{array}[]{c}\epsilon_{0,1}|{E_{0,1}}\rangle+\epsilon_{0,2}|{E_{0,2}}\rangle=0\\ \Downarrow\\ \epsilon_{0,1}=\epsilon_{0,2}=\epsilon_{0,3}=0.\end{array}$
@@
    (45) $\begin{array}[]{c}\underset{=0}{\epsilon_{0,2}}|{E_{0,2}}\rangle+\underset{=0}{\epsilon_{0,3}}|{E_{0,3}}\rangle=0.\end{array}$
@@
(46) $J_{{\operatorname{error}}}=J_{0}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{s_{0}}\rangle,|{s_{3}}\rangle,|{d_{-1}}\rangle,|{d_{0}}\rangle,|{d_{3}}\rangle\right\}.$
@@
    (47) $\epsilon_{1,-2}|{E_{1,-2}}\rangle-\epsilon_{1,-1}|{E_{1,-1}}\rangle=0.$
@@
    (48) $\begin{array}[]{c}\epsilon_{1,-1}|{E_{1,-1}}\rangle-\epsilon_{1,0}|{E_{1,0}}\rangle=0\\ \Downarrow\\ \epsilon_{1,-2}=\epsilon_{1,-1}=\epsilon_{1,0}.\end{array}$
@@
    (49) $\epsilon_{1,2}|{E_{1,2}}\rangle-\epsilon_{1,3}|{E_{1,3}}\rangle=0.$
@@
    (50) $\begin{array}[]{c}\epsilon_{1,-2}|{E_{1,-2}}\rangle+\epsilon_{1,-1}|{E_{1,-1}}\rangle=0\\ \Downarrow\\ \epsilon_{1,-2}=\epsilon_{1,-1}=\epsilon_{1,0}=0.\end{array}$
@@
    (51) $\underset{=0}{\epsilon_{1,-1}}|{E_{1,-1}}\rangle+\underset{=0}{\epsilon_{1,0}}|{E_{1,0}}\rangle=0.$
@@
    (52) $\begin{array}[]{c}\epsilon_{1,2}|{E_{1,2}}\rangle+\epsilon_{1,3}|{E_{1,3}}\rangle=0\\ \Downarrow\\ \epsilon_{1,2}=\epsilon_{1,3}=0.\end{array}$
@@
(54) $J_{{\operatorname{error}}}=J_{1}\cup J_{{\operatorname{invalid}}}=\left\{|{s_{-1}}\rangle,|{d_{-1}}\rangle,|{s_{1}}\rangle,|{s_{3}}\rangle,|{d_{3}}\rangle\right\}.$
@@
(55) $\begin{array}[]{c}\sum_{k}\epsilon_{0,k}\beta_{k,d_{1}}|{E_{0,k}}\rangle+\sum_{k}\epsilon_{1,k}\beta_{k,d_{1}}|{E_{1,k}}\rangle=0\\ \Downarrow\\ \frac{i}{\sqrt{2}}(-\epsilon_{0,0}|{E_{0,0}}\rangle+\underset{=0}{\epsilon_{0,1}}|{E_{0,1}}\rangle)+\frac{i}{\sqrt{2}}(-\underset{=0}{\epsilon_{1,0}}|{E_{1,0}}\rangle+\epsilon_{1,1}|{E_{1,1}}\rangle)=0\\ \Downarrow\\ \epsilon_{0,0}|{E_{0,0}}\rangle=\epsilon_{1,1}|{E_{1,1}}\rangle.\end{array}$
@@
(56) $\begin{array}[]{c}U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}},\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}\\ \Downarrow\\ U_{\mathrm{E}}|{0}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{A}}=|{\phi}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{A}}.\end{array}$